OpenStudy (anonymous):

ln(uv^3)^10=A ln u+B ln v what are A and B?

myininaya (myininaya):

\[\ln(ab)=\ln(a)+\ln(b)\] use this first

myininaya (myininaya):

on \[ln(uv^3)\]

myininaya (myininaya):

\[\ln(u)+\ln(v^3) \text{ like this}\]

myininaya (myininaya):

now use \[\ln(x^r)=r\ln(x)\] and you are basically done

OpenStudy (anonymous):

i dont get how i am suppose to enter these variables into my calc

myininaya (myininaya):

\[\ln(u)+3 \ln(v)\] I just used that rule I put

myininaya (myininaya):

\[1 \ln(u)+3 \ln(v)\]

myininaya (myininaya):

So now it is easy for you to determine which is A and which is B

myininaya (myininaya):

so what is A? and what is B?

myininaya (myininaya):

A ln(u) + B ln(v) 1 ln(u) + 3 ln(v) Do you see what A and B is?

OpenStudy (anonymous):

1 and 3?

myininaya (myininaya):

Oh I didn't see that ^10 at first

myininaya (myininaya):

\[\ln((uv^3)^{10})\] is that right?

OpenStudy (anonymous):

yea

myininaya (myininaya):

\[10 \ln(u v^3)=10 (\ln(u)+\ln(v^3))=10 \ln(u)+10 \ln(v^3)=10 \ln(u)+10(3) \ln(v)\]

myininaya (myininaya):

So A is? and B is?

OpenStudy (anonymous):

10 and 30 thanks for the help. i have 2 more similar but somewhat different

OpenStudy (anonymous):

Use the laws of logarithm to rewrite the expression log (x^14 y^12/ z^13) in a form with no logarithm of a product, quotient, or power. After rewriting we have log (x^14 y^12/ z^13)= A log(x) + B log(y) + C log(z). what are A, B, and C? i got a and b, just need c